### v1

# Giuseppe Scanniello+

Abstract
The management of chronic Heart Failure (HF) presents significant challenges in modern healthcare, requiring continuous monitoring, early detection of exacerbations, and personalized treatment strategies. In this paper, we present a predictive model founded on Machine Learning (ML) techniques to identify patients at HF risk. This model is an ensemble learning approach, a modified stacking technique, that uses two specialized models leveraging clinical and echocardiographic features and then a meta-model to combine the predictions of these two models. We initially assess the model on a real dataset and the obtained results suggest that it performs well in the stratification of patients at HR risk. Specifically, we obtained high sensitivity (95%), ensuring that nearly all high-risk patients are identified. As for accuracy, we obtained 84%, which can be considered moderate in some ML contexts. However, it is acceptable given our priority of identifying patients at risk of HF because they will be asked to participate in the telemonitoring program of the PrediHealth research project on which some of the authors of this paper are working. The initial findings also suggest that ML-based risk stratification models can serve as valuable decision-support tools not only in the PrediHealth project but also for healthcare professionals, aiding in early intervention and personalized patient management. To have a better understanding of the value and of potentiality of our predictive model, we also contrasted its results with those obtained by using three baseline models. The preliminary results indicate that our predictive model outperforms these baselines that flatly consider features, i.e., not grouping them in clinical and echocardiographic features.
Index Terms:
Healthcare, Machine Learning, Predictive Models, Artificial Intelligence, Digital Health
I Introduction
In modern healthcare systems, efficiently stratifying individuals [^1] It consists in categorizing them into different risk groups based on their health status, disease progression, or other factors. based on their risk profiles and predicted health trajectories is crucial for optimizing medical interventions, resource allocation, and personalized treatment strategies. The increasing availability of electronic health data, wearable sensor data, and advanced Machine Learning (ML) techniques has created new opportunities for predictive modeling in patient management. However, despite these advancements, significant challenges remain in ensuring the predictive models’ accuracy. In particular, the heterogeneity of patients and the complexity of disease progression demand stratification approaches that can adapt to diverse healthcare environments [1].
The human heart works like a pump, sending blood and oxygen around the body. A major cause of morbidity and mortality worldwide is chronic Heart Failure (HF), which is defined by the heart’s inability to pump blood efficiently. Lifestyle risk factors such as smoking, poor diet, and high blood pressure are closely associated with the increase in the incidence of heart disease. Current techniques often fail to provide prompt and precise identification of heart disease, even with advances in medical diagnostics [2, 3]. It is also worth mentioning that cardiovascular diseases are associated with the highest number of deaths among Non-Communicable Diseases, amounting to 17.9 million deaths annually, followed by cancers (9.3 million), chronic respiratory diseases (4.1 million), and diabetes (2 million, including kidney disease linked to diabetes) [4].
The PrediHealth project aims to address the challenges above by integrating telemedicine, mobile health (m-health) solutions, and predictive analytics into an IoT healthcare platform. Specifically, it leverages an interoperable IoT web-based application, a telemonitoring kit equipped with medical devices and environmental sensors, and ML/AI-driven predictive models to support clinical decisions. As it is easy to follow, at the core of the PrediHealth project lies the predictive model–designed for the stratification of patients at risk of HF–that is presented in this paper for the first time. This model is based on an ensemble learning approach–a modified stacking technique–that uses two specialized models leveraging clinical and echocardiographic features and then a meta-model to combine the predictions of these two models. In this paper, we detail the model’s methodological foundations and its initial empirical assessment, as well as the possible implications for our results outside the goal of the PrediHealth project. We assessed the validity of the model by conducting an initial empirical study. The obtained results suggest that the model performs pretty well. Specifically, we obtained high sensitivity (95%), ensuring that nearly all high-risk patients are identified. As for accuracy, we obtained 84%, which can be considered moderate in some ML contexts. However, it is acceptable in the PrediHealth project, given the priority of identifying patients at risk of HF to be then included in a telemonitoring program, to act promptly to let them receive medical attention, thus hopefully preventing them from dying. We also compared our model against three baseline models (Decision Tree, Random Forest, and Support Vector Regression). Our initial outcomes suggest that the use of our predictive model outperforms these baselines (e.g., in terms of the accuracy and the sensitivity of the patient stratification).
Paper Structure. The following sections will explore the model’s design and evaluation strategies, providing a perspective on its role in the future of predictive healthcare and in PrediHealth, in particular. Precisely, the rest of the paper is organized as follows. In Section II, we introduce the PrediHealth project and present related research and background information, while the design of the predictive model is shown in Section III. The design of the study to empirically assess our predictive model and the results of this assessment are reported in Section IV and Section V, respectively. In this latter section, we also discuss the results and delineate possible implications for our research. We conclude the paper in Section VI with final remarks and future work.
II Background and Related Work
In this section, we first provide information on the PrediHealth project. It is the project for which the predictive model presented in this paper has been developed. We conclude by motivating our work and discussing related research.
II-A The PrediHealth Project
The high-level workflow of PrediHealth begins with the stratification of individuals at risk of HF using AI/ML-powered predictive models. Individuals identified as patients at HF risk are then equipped with wearable and environmental sensors, which are seamlessly integrated into a telemonitoring system based on an IoT-enabled web platform. Accordingly, the primary objectives of the projects are:
O1: Develop an interoperable web-based IoT platform. The platform has to establish a secure, scalable, and standardized digital infrastructure for the seamless collection, transmission, and analysis of health data from telemonitoring devices. This platform facilitates real-time monitoring of HF patients by integrating wearable medical devices and environmental sensors, ensuring continuous and remote patient monitoring.
O2: Design a telemonitoring kit. This objective focuses on the development of a telemonitoring kit specifically designed for patients with HF, integrating advanced (wearable and not) medical devices and environmental sensors to enable continuous health monitoring.
O3: Develop a multimarker score to support decision-making. The goal is twofold: (i) building AI/ML-powered models to risk stratification of patients vulnerable to acute exacerbations or disease progression for HF and (ii) defining thresholds and scores for risk factors associated with HF exacerbations of patients when using telemonitoring kits (i.e., medical devices and environmental sensors identified in O2).
We restricted ourselves to a few aspects of the project for space reasons and scant relevance to the overarching goals of the research presented in the paper. Readers interested in further details can refer to .
II-B Predictive Models in Digital Health
Predictive models play a crucial role in the management of chronic diseases such as HF by enabling early detection, risk stratification, and personalized treatment for patients. These models analyze large datasets to support proactive interventions, reduce hospitalizations, and improve patient outcomes, making predictive modeling a key component of modern telemedicine and, more specifically, telemonitoring. Existing research has primarily focused on developing predictive models using electronic health records and demographic variables to enhance early detection and risk stratification. For example, Medhi et al. [5] conducted a review by selecting 150 studies from databases such as PubMed, Google Scholar, and the Cochrane Library. Their synthesis detailed how AI/ML-powered models—spanning classical methods like SVMs, Random Forest, and decision trees to advanced deep learning approaches—can process large datasets from wearables, imaging, and other clinical sources to enable early detection, risk stratification, and personalized treatment of heart failure. Their work highlights the potential of these predictive models to reduce hospitalizations and improve patient outcomes and emphasizes the transformative role of AI in modern telemedicine and telemonitoring.
Traditional models, such as trajectory-based disease progression models and multivariable Cox regression, have been used to identify key predictors of mortality and morbidity, including age, diabetes, and left ventricular ejection fraction [2]. These models often emphasize heart failure with reduced ejection fraction (HFrEF), leaving a gap in predictive capabilities for preserved ejection fraction (HFpEF) [3].
Other researchers have explored ML-driven models for HF prognosis, utilizing algorithms like logistic regression, decision trees, random forests, support vector machines (SVMs), and ensemble classifiers. For instance, Pinem et al. [6] demonstrated that a Voting Classifier combining Logistic Regression and SVM achieved the highest accuracy (88.04%) and ROC_AUC score (88%), with blood pressure and cholesterol emerging as key predictive factors. Similarly, Guo et al. [7] employed ML techniques to predict HF outcomes using EHR data, incorporating medical notes, laboratory results, and imaging data to achieve expert-level diagnostic accuracy. These studies underscore the role of ML in early HF detection, risk assessment, and personalized intervention strategies. Recently, Visco et al. [8] developed a predictive model for cardiovascular diagnostics. Their study identified correlations between worsening HF and three key parameters: Creatinine, sPAP, and CAD. To enhance interpretability, they introduced a GP-based classifier and extracted from it a formula that achieved almost 97% accuracy, outperforming traditional ML models. On the other hand, the authors themselves identified how the use of only four features, despite the high accuracy, could leave out from the doctor’s final evaluation other parameters that could be important in the final prediction.
Despite these advancements, existing HF risk models face notable challenges. Voors et al. [9] showed inconsistencies in multinational cohort data, which affect model generalizability, while Adler et al. [10] pointed out limitations in single-center studies, particularly the exclusion of elderly populations. Moreover, many ML models struggle with imbalanced datasets and lack interpretability, which hinders their adoption [11].
The research presented in this paper aims to bridge these gaps by developing a robust ML-based predictive model that integrates diverse patient datasets, including HFpEF cases, to enhance generalizability. The proposed model not only refines risk stratification but also enhances clinical applicability by ensuring interpretability and reliability across heterogeneous patient populations. In addition, our predictive model is integrated within advanced solutions in the context of CPS/IoT fields [^2] While CPS encompasses a broader range of networked systems, IoT specifically focuses on internet-enabled connectivity. by integrating state-of-the-art technologies such as telemonitoring and m-health to facilitate the integrated hospital-territory-home management of HF patients.
III Design of the Predictive Model
In this section, we first introduce the dataset employed for training and testing the model. We then describe the architecture of the proposed model in detail.
III-A Dataset
The used dataset contains real data. That is, it comprises clinical records of patients diagnosed with heart failure, collected as part of a structured medical registry. It encompasses a diverse set of 1040 patient records, each characterized by 66 features that provide a comprehensive view of patient demographics, clinical history, treatment details, and outcomes. The dataset includes key variables such as age, height, weight, and body mass index (BMI), along with clinical indicators like ejection fraction (EF), New York Heart Association (NYHA) classification, and specific diagnoses, including ischemic, hypertensive, and dilated cardiomyopathy. Additionally, dataset documents prescribed medications, including ACE inhibitors and aldosterone antagonists, offering insights into pharmacological management. Longitudinal follow-up data track patient progress, recording mortality outcomes such as cardiac death, heart failure exacerbation, sudden death, and interventions like heart transplants or ventricular assist device (VAD) implantation. The dataset also contains other bio-chemical pieces of information like Na+ (Sodium) or Uric Acid.
Before using the dataset, we preprocessed it to ensure data consistency and reliability. In other words, the dataset underwent pre-processing steps to address missing values, standardize feature formats, and encode categorical variables where necessary, ensuring consistency and reliability for subsequent modeling and evaluation. In the following sections, we detail the performed pre-processing steps and outline the criteria and methodology used for the selection of the features used in our predictive model, explaining the rationale behind the choice of these features and their relevance for the construction of such a model.
III-A1 Pre-processing
The dataset underwent the following pre-processing steps.
  • The gender encoding was standardized, as the dataset was originally formed by merging two datasets that used different encoding schemes. In one dataset, males were labeled as 1 and females as 0, whereas in the other, females were encoded as 1 and males as 2. To unify this, the gender values were mapped to a common scheme where male = 0 and female = 1.
  • All rows missing the target variable “death” were removed to maintain the integrity of the supervised learning process. Following this, rows containing more than three missing values were eliminated to reduce the impact of incomplete records on model performance.
  • Date values were converted into a standardized date-time format, and a new feature, “lifespan,” was introduced to represent the number of days between the initial characterization date and the recorded outcome date (“death”). Cases where the lifespan was negative were manually inspected and removed to ensure logical consistency in survival time calculations.
  • Outliers in the “lifespan” variable were identified and removed using the inter-quartile range (IQR) method. The lower and upper bounds were determined based on 1.5 times the IQR, and values outside this range were filtered out. The cleaned lifespan distribution was then analyzed to ensure a more balanced dataset.
  • As for the outcome variable (or also dependent variable), we used a binary classification for the values in the dataset. We employed a threshold-based labeling approach, using a 440-day threshold to classify patients into two groups: “at risk” and “not at risk.” Patients with a lifespan greater than or equal to 440 days were assigned label = 0, while those with a lifespan below this threshold were assigned label = 1. This transformation allowed for a more structured and interpretable analysis, enabling predictive modeling focused on early mortality risk assessment. The choice of this threshold is justified by technical reasons and clinical bases. Indeed, from a technical point of view, this threshold divides the dataset almost perfectly; in fact, we have 48% of samples with the label ”at risk” and 52% of samples with the label ”not at risk.” This also allowed the predictive model to exhibit good prediction performance of the outcome values. As far as the clinical base is concerned, the used threshold can be considered to be compliant with a period within which vigilant monitoring and proactive management are crucial to mitigate adverse outcomes in HF patients.
After completing all pre-processing steps shown just above, the final dataset consisted of 182 patient records to be used in our predictive model.
III-A2 Features
In our predictive model, we selected 33 out of 66 available features, including clinical and echocardiographic parameters relevant to cardiac evaluation. The chosen characteristics encompass demographic attributes such as age, BMI and sex, along with essential clinical indicators such as ejection fraction (EF), NYHA classification, NT-proBNP, creatinine, glucose, and hemoglobin (Hb). Additionally, the dataset contains information on comorbidities such as hypertension, diabetes, dyslipidemia, and respiratory diseases. The selection of features was based on the experience some of the authors gained in the field of HF, taking into account the purposes of the research presented in this paper and those of the PrediHealth project (e.g., stratifying individuals into patients at risk or not). Specifically, we decided to keep all data related to a standard clinical and instrumental cardiological examination (without null values). Therefore, key echocardiographic measurements were retained, including left ventricular mass index (LVMI), ventricular dimensions, atrial size, posterior wall thickness, septum thickness, right ventricular diameter, and tricuspid annular systolic excursion (TAPSE). Electrical abnormalities on electrocardiographic evaluation, such as left and right bundle branch blocks (BBSx, BBDx), atrial fibrillation (FA), atrial flutter, and pacemaker presence (PM), were also included. Height and weight were removed since they are inherently represented by the BMI feature. In contrast, other variables such as sodium, potassium, uric acid, urea, free triiodothyronine (FT3), and free thyroxine (FT4) were excluded as they were considered secondary to the main goal of our research. Some other features were deliberately excluded due to the high percentage of missing values. Given the medical nature of the dataset, we opted not to impute missing values with synthetic or inferred data to preserve the integrity of the analysis and avoid introducing potential biases.
Medication usage data, including beta-blockers, ACE inhibitors, and aldosterone antagonists, were incorporated to assess treatment effects. In addition, two derived characteristics were included: “lifespan,” representing the number of days between diagnosis and outcome, and “label,” a binary classification indicating whether a patient survived beyond the 440-day threshold. The features, grouped in “Clinical Features” and “Echocardiographic Features,” that serve as the foundation of our predictive model are detailed in Table I.
The feature categorization into Clinical and Echocardiographic Features reflects their complementary roles in HF assessment. Clinical Features capture systemic health, metabolic status, and hemodynamic regulation, which are essential for identifying HF risk factors. Key prognostic markers like NT-proBNP, creatinine, and NYHA classification provide insight into HF progression [12]. Echocardiographic Features offer direct evaluation of cardiac structure and function. Metrics such as EF, LVMI, and TAPSE help distinguish HF phenotypes and guide treatment. The inclusion of ECG abnormalities like bundle branch blocks and atrial fibrillation enhances prognosis by detecting conduction disturbances. This grouping integrates systemic and cardiac indicators, improving risk stratification.
TABLE I: Description of the features used in the dataset.
Feature Description Type
Clinical Features
Primary_Diagnosis Primary diagnosis of the patient Categorical
Secondary_Diagnosis Divided into Primary and Secondary if two diagnoses are present Categorical
HFpEF Heart failure with preserved ejection fraction classification Categorical
EF Ejection Fraction (%) – Left ventricular pumping function Numeric
NYHA New York Heart Association (NYHA) class (I-IV), measures heart failure severity Numeric
Age Age of patients in years Numeric
BMI Body Mass Index Numeric
Sex Sex of patients Categorical (1 = Male, 0 = Female)
Hypertension High blood pressure Categorical (1 = Yes, 0 = No)
Dyslipidemia Abnormal lipid levels Categorical (1 = Yes, 0 = No)
Diabetic Blood sugar disorder Categorical (1 = Yes, 0 = No)
Bronchopneumonia Lung disease Categorical (1 = Yes, 0 = No)
Beta-Blocker Use of beta-blockers Categorical (1 = Yes, 0 = No)
ACE_SART Use of ACE inhibitors or sartans Categorical (1 = Yes, 0 = No)
Anti-Aldosterone Use of aldosterone antagonists Categorical (1 = Yes, 0 = No)
Echocardiographic Features
PARETE POST Posterior heart wall thickness Numeric
SETTO Septum thickness Numeric
LVES_DIAM Left ventricular end-systolic diameter Numeric
LVED_DIAM Left ventricular end-diastolic diameter Numeric
VDx (PARAST) Right ventricular diameter Numeric
LVMI Left ventricular mass index Numeric
ASx Left atrium size Numeric
TAPSE Tricuspid annular systolic excursion Numeric
RS Respiratory syndrome Categorical (1 = Yes, 0 = No)
BBSx Left bundle branch block Categorical (1 = Yes, 0 = No)
BBDx Right bundle branch block Categorical (1 = Yes, 0 = No)
NT-proBNP Heart failure biomarker Numeric
Blood Creatinine Level Kidney function marker Numeric
Glucose Blood sugar level Numeric
FA Atrial fibrillation Categorical (1 = Yes, 0 = No)
Flutter Atrial flutter Categorical (1 = Yes, 0 = No)
PM Pacemaker presence Categorical (1 = Yes, 0 = No)
Hb Hemoglobin level Numeric
III-B Predictive Model Workflow
Our predictive model is based on an ensemble learning approach, a modified stacking technique. Standard stacking combines multiple heterogeneous specialized models trained on the same feature set, our model employs domain-specific feature segregation to enhance predictive accuracy. Specifically, the input data are divided into two distinct feature groups: clinical features, including demographic data, patient history, laboratory results, and risk factors, and echocardiographic features. Each feature group is processed separately using two identical ML models, each trained exclusively on one of the feature sets. This separation ensures that each model specializes in learning patterns from its respective data type without interference from other features. In our implementation, both the specialized models and the meta-model are constructed using Random Forest algorithms. In detail, each specialized model is tuned using a five-fold cross-validation strategy, where the training data are partitioned into five subsets. In detail, for each specialized model, we employ a five-fold cross-validation strategy using GridSearchCV, which systematically tests various combinations of hyperparameters, such as the number of trees, maximum depth, and minimum samples required for a split, to determine the best configuration that minimizes prediction error. Once these models are optimized, a single train-test split (with 20% of the data reserved for testing) is utilized to evaluate the final performance. The outputs of the two optimized specialized models on the test set are then used as input features for the meta-model. Our predictive model workflow is inspired by the research presented by Pinem et al. [6], where the authors trained multiple models on the same set of features to predict heart failure risk. In that work, the authors used the output of the models as input for a Voting Classifier. In our case, the meta-model on top is also trained on the output of the specialized models, but the main difference is that each specialized model is trained using different features.
The outcome variable of the two models (i.e., the output of the two specialized models) is the estimated lifespan of a patient. The outputs of the two specialized models represent the input for a meta-model, which integrates their predictions to generate the final category for the patients. Specifically, the final output of the meta-model is expressed in terms of estimated lifespan. As mentioned before, the individuals for whom the model predicts a lifespan of less than 440 days are considered as patients “at risk,” while the other are considered as “not at risk.” In the context of the PrediHealt project, the identification of patients at HF risk allows them to participate in a telemonitoring program, with the objective of acting promptly to let them receive medical attention.
Integrating two specialized models within a metamodel offers several potential advantages over a single ML model trained on all features. Specifically, these advantages are:
  1. 1.
    Improved Interpretability and Modularity. By separating clinical and echocardiographic features, the contribution of each data type can be analyzed independently. This modular approach allows for better insight into how different variables influence predictions.
  2. 2.
    Reduced Risk of Overfitting. A single model trained on all features may overfit due to redundant or spurious correlations. By using specialized models for each group of features (clinical and echocardiographic), we reduce the risk of one type of feature dominating the learning process, improving generalization.
  3. 3.
    Increased Robustness. This structure makes the system less sensitive to missing data. If a patient lacks echocardiographic measurements, the clinical model can still contribute to predictions, preventing the complete failure of the model. The metamodel also balances predictions, reducing the impact of errors from individual models.
  4. 4.
    Data availability. We built a model based on variables readily available in clinical reality. All the components used represent part of a standard cardiac evaluation [13].
  5. 5.
    Greater Adaptability and Scalability. If new predictive techniques emerge for clinical or echocardiographic data, only the respective model needs to be updated, rather than retraining an entire model on all the features. This characteristic allows for continuous improvement of the predictions while maintaining the same structure of the predictive framework.
IV Study Design
In this section, we show the design of our empirical study. We took care of the suggestions/guidelines by Wohlin et al. [14] and Juristo and Moreno [15] to design and conduct experiments in the software engineering field.
IV-A Research Question
The primary objective of the experimental assessment is to assess preliminarily the validity of the developed predictive model for the stratification of patients at HR risk. In particular, we aimed to show whether the proposed predictive model, which is based on two specialized models leveraging clinical and echocardiographic features and then a meta-model to combine the predictions of these two models, outperforms traditional ML models trained on the entire feature set (i.e., the features used in the two specialized models). To that end, we considered three baseline models–i.e., Decision Tree, Random Forest, and Support Vector Regression (SVR)– and compared their results (e.g., accuracy, precision, and sensitivity) with those of our prediction model. We choose these baselines because they represent the leading ML techniques in several AI application contexts and medical domains [16], in particular.
  • Decision Tree. It is a tree-structured model that recursively splits the data based on feature values to predict an outcome. It is intuitive and interpretable and can handle both classification and regression tasks.
  • Random Forest. It is an ensemble learning method that builds multiple decision trees on random subsets of data and features and aggregates their predictions. It improves generalization by reducing overfitting (e.g., as compared with the Decision tree) and increasing robustness. It is well-suited for handling high-dimensional data and capturing complex patterns.
  • Support Vector Regression. It is a regression variant of Support Vector Machines that aims to find a function within a specified error margin (epsilon) while maximizing model flatness. It captures non-linear relationships using kernel functions.
For the baseline models, as we did for our predictive model, the individuals are arranged into the two groups according to the lifespan prediction, i.e., if it is less than 440 days, the individual is considered at risk, otherwise not.
According to the primary objective of our experimental assessment, we defined and investigated the following main Research Question:
RQ: Does our predictive model outperform leading ML techniques in classifying at-risk patients?
IV-B Experimental Methodology
To study RQ, we conducted a comparative analysis between our model and the three baselines mentioned before. These baselines, differently from our model, were trained using the entire set of 33 selected features in Table I. All models were trained and evaluated using the final dataset, consisting of 182 patient records, with 20% of the data reserved for testing. Baselines were implemented using standard algorithms from the scikit-learn library in Python. Just as for training the models underlying our meta-model, the baseline models have been optimized using a five-fold cross-validation strategy and GridSearchCV, which systematically tests various combinations of hyperparameters. This methodology ensures that all models were evaluated using the same data splits, enabling a fair comparison when studying RQ.
IV-C Metrics to Assess the Predictions of the Models
The evaluation metrics used are standard and widely accepted in the application of ML techniques within the medical domain [17]. The used metrics are:
  • Accuracy: It represents the proportion of patients that are correctly identified and it is computed as follows:
    $\text{Accuracy}=\frac{TP+TN}{TP+TN+FP+FN}\%$
    This metric assumes values between 0% and 100%. The higher the value, the better it is in terms of the accuracy of the prediction. For the sake of clarity and completeness, TPs (True Positives) refer to patients correctly identified as people at HF risk, while TNs (True Negatives) refer to patients correctly identified as patients not at HF risk. Instead, FPs (False Positives) represent patients mistakenly identified as at risk, while FNs (False Negatives) represent patients mistakenly identified as not at risk.
  • Precision (for the at-risk class): It indicates the proportion of patients predicted as at risk who truly are at risk. It is calculated as follows:
    $\text{Precision}=\frac{TP}{TP+FP}\%$
    Also, the precision metric assumes values between 0% and 100%. The higher the value, the better the correctness of the prediction is.
  • Sensitivity (Recall for the at-risk class): It measures the model’s ability to correctly identify patients at risk and is defined as follows:
    $\text{Sensitivity}=\frac{TP}{TP+FN}\%$
    Also, this metric assumes values between 0% and 100%. The higher the value, the better the sensitivity of the model is.
  • F1-Score (for the at-risk class): It is the balanced harmonic mean of precision and sensitivity. It offers a balanced measure when the costs of false positives and false negatives differ and is computed as follows:
    $F1\text{-Score}=2\times\frac{\text{Precision}\times\text{Sensitivity}}{\text{Precision}+\text{Sensitivity}}\%$
    This metric assumes values between 0% and 100%. The higher the value, the better it is.
  • Diagnostic Odds Ratio (DOR): It is a comprehensive metric that integrates sensitivity and specificity. It is computed as follows:
    $DOR=\frac{TP\times TN}{FP\times FN}$
    The higher the value for DOR, the better it is.
It is important to highlight that, since the model is intended to stratify patients by identifying those at risk for inclusion in the telemonitoring program, minimizing FN is a primary objective. In this context, high precision is particularly desirable, as it ensures that most of the patients flagged as “at risk” truly need intervention. Nevertheless, achieving high values across all evaluation metrics remains a desirable goal to ensure overall model reliability.
TABLE II: Medical evaluation metrics of the proposed model
Metric Value Explanation of the Results
Accuracy 84% 84% of patients are correctly identified as at risk.
Precision (At-Risk) 79% 79% of patients flagged as at risk truly are at risk.
Sensitivity (At-Risk) 95% 95% of at-risk patients are correctly identified.
F1-Score (At-Risk) 86% Good balance between precision and sensitivity.
DOR 45.6 Strong discriminative ability.
V Results
In this section, we highlight the main findings. We conclude by showing possible practical implications and threats that could affect the validity of the obtained results.
V-A Main Findings and Discussion
The performance results of our predictive model are summarized in Table II. While the overall accuracy of 84% may appear moderate at first glance, the high sensitivity (95%) ensures that nearly all high-risk patients are correctly identified—an essential requirement for patient monitoring in the PrediHealth project. The precision of 79% reflects a relatively high number of false positives, which is acceptable given the clinical context and the overall objective of the PrediHealth project by minimizing FNs. Furthermore, the high Diagnostic Odds Ratio (45.6) underscores the model’s strong discriminative power between high- and low-risk patients, reinforcing its suitability for deployment within PrediHealth.
The results in Table III suggest that the proposed model significantly outperforms the baseline machine learning models. It achieves an accuracy of 84%, compared to 57% for the Decision Tree, 60% for the Random Forest, and 62% for SVR. Notably, the sensitivity of our model reaches 95% which is well above the 60% (Decision Tree and Random Forest) and 50% (SVR) by suggesting a strong ability to correctly identify high-risk patients as compared with the baseline models. The precision of our model (79%) is slightly higher than that of SVR (71%), which has the highest precision among the baseline models. This indicates that our model has a better balance between minimizing false positives and ensuring the inclusion of at-risk patients in the telemonitoring program.
Additionally, the proposed model achieves an F1-score of 86%, outperforming Decision Tree 60% and Random Forest (62%), and Support Vector Regression (59%). Its DOR (45.6) further confirms its superior ability to distinguish between high-risk and low-risk patients, compared to Decision Tree (1.69), Random Forest (2.14), and SVR (3.25). These improvements highlight the model’s reliability for HF risk stratification of patients.
TABLE III: Medical evaluation metrics of the baseline ML models
ML Model Metric Value
Decision Tree Accuracy 57%
Precision (At-Risk) 60%
Sensitivity (At-Risk) 60%
F1-Score (At-Risk) 60%
DOR 1.69
Random Forest Accuracy 60%
Precision (At-Risk) 63%
Sensitivity (At-Risk) 60%
F1-Score (At-Risk) 62 %
DOR 2.14
Support Vector Regression Accuracy 62%
Precision (At-Risk) 71%
Sensitivity (At-Risk) 50%
F1-Score (At-Risk) 59%
DOR 3.25
V-B Implications
Here, we focus on the implications of our results from the perspectives of practitioners and researchers, setting aside their specific relevance to the PrediHealth project. This choice reflects our view that the project’s implications are already evident, and it allows us to emphasize that our findings hold significance beyond the boundaries of the project itself. Specifically, we developed a predictive model for HF risk stratification of patients using a modified stacking ensemble approach applied to clinical and echocardiographic data. It achieved high sensitivity while maintaining an adequate accuracy, using a 440-day lifespan threshold. This outcome suggests that a modified stacking ensemble approach can effectively identify a significant portion of high-risk HF patients. This is relevant for practitioners in the healthcare field, as the model functions as a valuable screening tool. Its high sensitivity supports its role in initial patient stratification, prompting closer monitoring. Researchers could be interested in improving our model in order to enhance its accuracy while maintaining high sensitivity, potentially by integrating additional data modalities or optimizing our stacking ensemble approach.
Our overall model, which separates clinical and echocardiographic features into distinct specialized models before integrating their predictions via a meta-model, presents a potential advantage, as our initial results suggest. Similar to how modularity improves software systems, this structure may enhance the interpretability of risk scores for clinicians. Practitioners could benefit from this increased transparency, as it allows them to discern whether a patient’s risk is primarily driven by general clinical status or specific cardiac parameters derived from echocardiography. Researchers should empirically validate the interpretability and robustness of this architecture, comparing it to monolithic models trained on all features flatly. Our initial results seem to justify this future research direction and also to explore optimal meta-learning strategies for integrating predictions from specialized models by assessing the scalability when incorporating additional features.
The predictive model was trained and evaluated on real patient data/records, which underwent extensive pre-processing, including feature selection and the removal of records with missing data, ultimately resulting in a dataset of 182 instances. Researchers could validate the model’s performance and generalizability on larger, multi-center datasets representing diverse patient populations. A particular focus on HFpEF cases, as discussed in our related work, would be valuable. Researchers could also be interested in investigating the impact of different missing data imputation techniques and the development of automated, context-aware feature selection methods to enhance the model’s robustness and reliability.
V-C Threats to Validity
In the following section, we discuss threats that could affect the validity of our study.
V-C1 External validity
Our predictive model was developed and evaluated using 182 samples. Generalizing outcomes from this clinically specific cohort poses a threat to external validity. The model’s performances might differ in other populations with varying demographic profiles, healthcare practices, or different diagnoses. Furthermore, the specific set of 33 selected features might not be universally collected or hold the same predictive power in different clinical contexts. Although our study provides initial evidence for the feasibility of this model within the PrediHealth project, caution is needed when generalizing our outcomes.
V-C2 Internal Validity
Internal validity threats mostly concern factors internal to our study design that might have influenced the results. The main internal validity threat of the study concerns the limited number of baselines used for comparison. This is because this does not exclude that there could be other ML models with better performance. On the other hand, the models used as comparison baselines are among the most used in the medical field, justifying their choice.
V-C3 Construct Validity
As for this kind of threats to validity, one key concern is how we define “HF risk” as survival of fewer than 440 days after characterization. While this provides a clear target, it may not fully capture the broader clinical view of when a patient needs intervention or faces a high risk of worsening health, which can lead to serious complications beyond death. As for the metrics used to assess our predictive model and baselines, we used standard performance measures such as sensitivity and accuracy. While these provide useful information, they possibly do not give a complete picture. Lastly, the model assumes that medical records are accurate. Any errors in these records could reduce the reliability of predictions and limit its usefulness in real-world settings.
V-C4 Conclusion validity
Our definition of the binary outcome (“at risk” vs. “not at risk”), chosen for technical reasons (dataset balance) and clinical relevance, could have affected outcomes. Using a different threshold could lead to different conclusions. Another internal validity threat concerns the choice of 33 features, which was based on clinical experience and data availability. This selection might leave out important predictors or include less useful ones, making it harder to fully represent the complex nature of HF risk.
V-C5 Reliability validity
Reliability validity threats concern the possibility of replicating our study. Due to the sensitive nature of patient health information and regulations like GDPR, the raw patient-level dataset used in this study cannot be made publicly available. This inherently limits direct replication. To mitigate this, we have detailed the features used in the model and the design of the empirical study to assess its validity.
VI Conclusions
In this paper, we propose a predictive model based on machine learning (ML) techniques to identify patients at risk of HF. The model adopts an ensemble learning approach using a modified stacking technique, which combines the outputs of two specialized models—one based on clinical features and the other on echocardiographic features—via a meta-model. We evaluate the model on a real-world dataset, and the results demonstrate strong performance in stratifying patients by HF risk. In particular, the model achieves high sensitivity (95%), ensuring that nearly all high-risk patients are correctly identified. The overall accuracy is 84%, which, while moderate in general ML contexts, is acceptable given our priority of minimizing false negatives (patients mistakenly identified as at risk). This is especially relevant for the telemonitoring program of the PrediHealth research project. Indeed, our preliminary findings also suggest that ML-based risk stratification models can serve as valuable decision-support tools—not only within the PrediHealth initiative but also more broadly in clinical practice—to facilitate early intervention and personalized patient management. We also compared our model against three baselines that treat features in a flat manner (i.e., without grouping them into clinical and echocardiographic categories). The results indicate that our model consistently outperforms these baselines, highlighting the benefit of feature specialization and structured integration.
Future research will focus on expanding the dataset to further validate the model, optimizing its interpretability for clinical applications, and integrating additional physiological and behavioral parameters to improve prediction. It is also worth noting that the current model was trained on real-world patient data from the Tuscany region of Italy. Due to the limited dataset size and its regional scope, the findings should be interpreted with caution; however, the promising initial results motivate conducting broader studies.
Acknowledgment
This project has been financially supported by the European Union NEXTGenerationEU project and by the Italian Ministry of the University and Research MUR (M4C2) as part of the project “Tuscany Health Ecosystem” - THE - Spoke 10 - CUP - J13C22000420001.

### v3

@@
-Giuseppe Scanniello+
+Machine Learning Solutions Integrated in an IoT Healthcare Platform for Heart Failure Risk Stratification
 Abstract
-The management of chronic Heart Failure (HF) presents significant challenges in modern healthcare, requiring continuous monitoring, early detection of exacerbations, and personalized treatment strategies.
-In this paper, we present a predictive model founded on Machine Learning (ML) techniques to identify patients at HF risk.
-This model is an ensemble learning approach, a modified stacking technique, that uses two specialized models leveraging clinical and echocardiographic features and then a meta-model to combine the predictions of these two models.
-We initially assess the model on a real dataset and the obtained results suggest that it performs well in the stratification of patients at HR risk.
-Specifically, we obtained high sensitivity (95%), ensuring that nearly all high-risk patients are identified.
-As for accuracy, we obtained 84%, which can be considered moderate in some ML contexts.
-However, it is acceptable given our priority of identifying patients at risk of HF because they will be asked to participate in the telemonitoring program of the PrediHealth research project on which some of the authors of this paper are working.
-The initial findings also suggest that ML-based risk stratification models can serve as valuable decision-support tools not only in the PrediHealth project but also for healthcare professionals, aiding in early intervention and personalized patient management.
-To have a better understanding of the value and of potentiality of our predictive model, we also contrasted its results with those obtained by using three baseline models.
-The preliminary results indicate that our predictive model outperforms these baselines that flatly consider features, i.e., not grouping them in clinical and echocardiographic features.
+This paper presents a predictive model founded on Machine Learning (ML) techniques to identify patients at Heart Failure (HF) risk.
+This model is an ensemble learning approach, a modified stacking technique, that uses three ML models leveraging clinical and echocardiographic features and then a meta-model to combine the predictions of these three models.
+We assessed the model on a real dataset, and our results suggest that it performs well in the stratification of patients at HF risk.
+Specifically, we obtained high sensitivity (91%) and good accuracy (78%).
+To better study the value and potentiality of our model, we also contrasted its results with those obtained by using several baseline ML models.
+The preliminary results indicate that our predictive model outperforms these baselines that flatly consider features, i.e., not grouping features in clinical and echocardiographic features.
+The proposed model is a part of the PrediHealth research project.
+This project integrates telemedicine, mobile health solutions, and predictive analytics into an IoT healthcare platform to manage patients at HF risk.
 Index Terms:
@@
 A major cause of morbidity and mortality worldwide is chronic Heart Failure (HF), which is defined by the heart’s inability to pump blood efficiently.
-Lifestyle risk factors such as smoking, poor diet, and high blood pressure are closely associated with the increase in the incidence of heart disease.
+Ageing of population and lifestyle risk factors such as smoking, poor diet, and high blood pressure are closely associated with the increase in the incidence of heart disease.
 Current techniques often fail to provide prompt and precise identification of heart disease, even with advances in medical diagnostics [2, 3].
@@
 The PrediHealth project aims to address the challenges above by integrating telemedicine, mobile health (m-health) solutions, and predictive analytics into an IoT healthcare platform.
-Specifically, it leverages an interoperable IoT web-based application, a telemonitoring kit equipped with medical devices and environmental sensors, and ML/AI-driven predictive models to support clinical decisions.
-As it is easy to follow, at the core of the PrediHealth project lies the predictive model–designed for the stratification of patients at risk of HF–that is presented in this paper for the first time.
-This model is based on an ensemble learning approach–a modified stacking technique–that uses two specialized models leveraging clinical and echocardiographic features and then a meta-model to combine the predictions of these two models.
-In this paper, we detail the model’s methodological foundations and its initial empirical assessment, as well as the possible implications for our results outside the goal of the PrediHealth project.
+Specifically, it leverages an interoperable IoT web-based application, a telemonitoring kit equipped with medical devices and environmental sensors, and ML/AI-driven predictive models to support clinical decisions [5].
+As it is easy to follow, at the core of the PrediHealth project lies the predictive model–designed for the stratification of patients at HF risk–that is presented in this paper for the first time.
+This model is based on an ensemble learning approach–a modified stacking technique–that uses three ML models leveraging clinical and echocardiographic features and then a meta-model to combine the predictions of these three bottom models.
+In this paper, we detail our model and its initial empirical assessment, as well as the possible implications for our results inside and outside the goal of the PrediHealth project.
 We assessed the validity of the model by conducting an initial empirical study.
 The obtained results suggest that the model performs pretty well.
-Specifically, we obtained high sensitivity (95%), ensuring that nearly all high-risk patients are identified.
-As for accuracy, we obtained 84%, which can be considered moderate in some ML contexts.
-However, it is acceptable in the PrediHealth project, given the priority of identifying patients at risk of HF to be then included in a telemonitoring program, to act promptly to let them receive medical attention, thus hopefully preventing them from dying.
-We also compared our model against three baseline models (Decision Tree, Random Forest, and Support Vector Regression).
-Our initial outcomes suggest that the use of our predictive model outperforms these baselines (e.g., in terms of the accuracy and the sensitivity of the patient stratification).
+We obtained high sensitivity (91%), ensuring that such a model nearly identifies all HF patients.
+As for accuracy, we obtained 78%.
+It is acceptable in the PrediHealth project, given the priority of identifying patients at risk of HF to be then included in a telemonitoring program, to act promptly to let them receive medical attention by using the IoT healthcare platform developed in the project.
+We also compared our model against three baseline models (e.g., Decision Tree and Support Vector Regression).
+Our initial outcomes suggest that the use of our model outperforms the baseline models (e.g., in terms of the accuracy and the sensitivity).
 Paper Structure.
-The following sections will explore the model’s design and evaluation strategies, providing a perspective on its role in the future of predictive healthcare and in PrediHealth, in particular.
-Precisely, the rest of the paper is organized as follows.
 In Section II, we introduce the PrediHealth project and present related research and background information, while the design of the predictive model is shown in Section III.
@@
 It is the project for which the predictive model presented in this paper has been developed.
-We conclude by motivating our work and discussing related research.
+Then we continue motivating our work and discussing related research.
+We conclude with a brief introduction to the ML models used in our work.
 II-A The PrediHealth Project
@@
 We restricted ourselves to a few aspects of the project for space reasons and scant relevance to the overarching goals of the research presented in the paper.
-Readers interested in further details can refer to .
+Readers interested in further details can refer to [5].
 II-B Predictive Models in Digital Health
@@
 [9] developed a predictive model for cardiovascular diagnostics.
-Their study identified correlations between worsening HF and three key parameters: Creatinine, sPAP, and CAD.
+Their study identified correlations between worsening HF and three key parameters: Creatinine, systolic pulmonary artery pressure (sPAP), and coronary artery disease(CAD).
 To enhance interpretability, they introduced a GP-based classifier and extracted from it a formula that achieved almost 97% accuracy, outperforming traditional ML models.
-On the other hand, the authors themselves identified how the use of only four features, despite the high accuracy, could leave out from the doctor’s final evaluation other parameters that could be important in the final prediction.
+However, the authors themselves identified how the use of only four features, despite the high accuracy, could leave out other parameters that could be important in the final prediction from the doctor’s final evaluation.
 Despite these advancements, existing HF risk models face notable challenges.
@@
 Moreover, many ML models struggle with imbalanced datasets and lack interpretability, which hinders their adoption [12].
-The research presented in this paper aims to bridge these gaps by developing a robust ML-based predictive model that integrates diverse patient datasets, including HFpEF cases, to enhance generalizability.
+The research presented in this paper aims to bridge these gaps by developing an ML-based predictive model that integrates diverse patient datasets, including HFpEF cases, to enhance generalizability.
 The proposed model not only refines risk stratification but also enhances clinical applicability by ensuring interpretability and reliability across heterogeneous patient populations.
@@
 by integrating state-of-the-art technologies such as telemonitoring and m-health to facilitate the integrated hospital-territory-home management of HF patients.
+Finally, integrating three models within a metamodel offers several advantages over a single ML model trained on all features.
+We detail these advantages in Section III-B.
+II-C Used ML-Models
+We used two machine learning models: the Logistic Regression Classifier and the Random Forest Classifier.
+The former is a linear model widely applied in binary classification problems [13].
+It applies a weighted sum of the input features and uses the logistic function to estimate the probability of a given class.
+Due to its simplicity, it is computationally efficient and generally less prone to overfitting, especially in low-dimensional spaces.
+These characteristics, along with the inherent interpretability of the model’s results (i.e., how easily a person can trace and explain the output based on the input features), make the Logistic Regression Classifier particularly valuable in clinical settings [14], where understanding model decisions is of central importance.
+In contrast, the Random Forest Classifier is a non-linear, ensemble-based method that constructs a collection of decision trees during the training phase and aggregates their outputs to form a final prediction [15].
+This approach enables the model to capture intricate patterns and interactions across features, particularly in heterogeneous or high-dimensional datasets.
+While Random Forests are generally robust and versatile, they tend to lack the transparency of linear models, which may be a drawback in domains where interpretability is a key requirement.
 III Design of the Predictive Model
-In this section, we first introduce the dataset employed for training and testing the model. We then describe the architecture of the proposed model in detail.
+In this section, we first introduce the dataset employed for the training and the testing of our model, and then we describe it in detail.
 III-A Dataset
-The used dataset contains real data.
+The used dataset contains real medical data.
 That is, it comprises clinical records of patients diagnosed with heart failure, collected as part of a structured medical registry.
-It encompasses a diverse set of 1040 patient records, each characterized by 66 features that provide a comprehensive view of patient demographics, clinical history, treatment details, and outcomes.
+It encompasses a diverse set of 1040 patient records, each characterized by unique ID and 65 features that provide a comprehensive view of patient demographics, clinical history, treatment details, and outcomes.
 The dataset includes key variables such as age, height, weight, and body mass index (BMI), along with clinical indicators like ejection fraction (EF), New York Heart Association (NYHA) classification, and specific diagnoses, including ischemic, hypertensive, and dilated cardiomyopathy.
-Additionally, dataset documents prescribed medications, including ACE inhibitors and aldosterone antagonists, offering insights into pharmacological management.
-Longitudinal follow-up data track patient progress, recording mortality outcomes such as cardiac death, heart failure exacerbation, sudden death, and interventions like heart transplants or ventricular assist device (VAD) implantation.
-The dataset also contains other bio-chemical pieces of information like Na+ (Sodium) or Uric Acid.
-Before using the dataset, we preprocessed it to ensure data consistency and reliability.
-In other words, the dataset underwent pre-processing steps to address missing values, standardize feature formats, and encode categorical variables where necessary, ensuring consistency and reliability for subsequent modeling and evaluation.
-In the following sections, we detail the performed pre-processing steps and outline the criteria and methodology used for the selection of the features used in our predictive model, explaining the rationale behind the choice of these features and their relevance for the construction of such a model.
-III-A1 Pre-processing
-The dataset underwent the following pre-processing steps.
-  • The gender encoding was standardized, as the dataset was originally formed by merging two datasets that used different encoding schemes.
-In one dataset, males were labeled as 1 and females as 0, whereas in the other, females were encoded as 1 and males as 2.
-To unify this, the gender values were mapped to a common scheme where male = 0 and female = 1.
-  • All rows missing the target variable “death” were removed to maintain the integrity of the supervised learning process.
-Following this, rows containing more than three missing values were eliminated to reduce the impact of incomplete records on model performance.
-  • Date values were converted into a standardized date-time format, and a new feature, “lifespan,” was introduced to represent the number of days between the initial characterization date and the recorded outcome date (“death”).
-Cases where the lifespan was negative were manually inspected and removed to ensure logical consistency in survival time calculations.
-  • Outliers in the “lifespan” variable were identified and removed using the inter-quartile range (IQR) method.
-The lower and upper bounds were determined based on 1.5 times the IQR, and values outside this range were filtered out.
-The cleaned lifespan distribution was then analyzed to ensure a more balanced dataset.
-  • As for the outcome variable (or also dependent variable), we used a binary classification for the values in the dataset.
-We employed a threshold-based labeling approach, using a 440-day threshold to classify patients into two groups: “at risk” and “not at risk.” Patients with a lifespan greater than or equal to 440 days were assigned label = 0, while those with a lifespan below this threshold were assigned label = 1.
-This transformation allowed for a more structured and interpretable analysis, enabling predictive modeling focused on early mortality risk assessment.
-The choice of this threshold is justified by technical reasons and clinical bases.
-Indeed, from a technical point of view, this threshold divides the dataset almost perfectly; in fact, we have 48% of samples with the label ”at risk” and 52% of samples with the label ”not at risk.” This also allowed the predictive model to exhibit good prediction performance of the outcome values.
-As far as the clinical base is concerned, the used threshold can be considered to be compliant with a period within which vigilant monitoring and proactive management are crucial to mitigate adverse outcomes in HF patients.
-After completing all pre-processing steps shown just above, the final dataset consisted of 182 patient records to be used in our predictive model.
-III-A2 Features
-In our predictive model, we selected 33 out of 66 available features, including clinical and echocardiographic parameters relevant to cardiac evaluation.
-The chosen characteristics encompass demographic attributes such as age, BMI and sex, along with essential clinical indicators such as ejection fraction (EF), NYHA classification, NT-proBNP, creatinine, glucose, and hemoglobin (Hb).
+Additionally, dataset documents prescribed medications, including ACE inhibitors, Beta-blockers, and aldosterone antagonists, offering insights into pharmacological management.
+Longitudinal follow-up data track patient progress, recording mortality outcomes such as death, heart failure exacerbation, and interventions like heart transplants or ventricular assist device (VAD) implantation.
+The dataset also contains, e.g., other bio-chemical pieces of information like Na+ (Sodium) or Uric Acid.
+III-A1 Features
+We extracted 33 out of 65 available features in the dataset, including clinical and echocardiographic parameters relevant to cardiac evaluation based on the experience of some of the authors in the HF field and our knowledge of the state of the art (e.g., [9]).
+The chosen characteristics encompass demographic attributes such as age, BMI, and sex, along with essential clinical indicators such as ejection fraction (EF), NYHA classification, NT-proBNP, creatinine, glucose, and hemoglobin (Hb).
 Additionally, the dataset contains information on comorbidities such as hypertension, diabetes, dyslipidemia, and respiratory diseases.
@@
 In contrast, other variables such as sodium, potassium, uric acid, urea, free triiodothyronine (FT3), and free thyroxine (FT4) were excluded as they were considered secondary to the main goal of our research.
+It is important to note that the Diagnosis feature was divided into two categorical features to enhance the granularity of the information and allow for a clearer differentiation between the identified diagnoses.
 Some other features were deliberately excluded due to the high percentage of missing values.
@@
 Medication usage data, including beta-blockers, ACE inhibitors, and aldosterone antagonists, were incorporated to assess treatment effects.
-In addition, two derived characteristics were included: “lifespan,” representing the number of days between diagnosis and outcome, and “label,” a binary classification indicating whether a patient survived beyond the 440-day threshold.
-The features, grouped in “Clinical Features” and “Echocardiographic Features,” that serve as the foundation of our predictive model are detailed in Table I.
+In addition, two derived characteristics were included: “lifespan,” representing the number of days between diagnosis and outcome, and “label.” The features, grouped in “Clinical Features” and “Echocardiographic Features,” that serve as the foundation of our predictive model are detailed in Table I.
 The feature categorization into Clinical and Echocardiographic Features reflects their complementary roles in HF assessment.
@@
 EF Ejection Fraction (%) – Left ventricular pumping function Numeric
-NYHA New York Heart Association (NYHA) class (I-IV), measures heart failure severity Numeric
+NYHA New York Heart Association (NYHA) class (I-IV), measures heart failure severity Categorical
 Age Age of patients in years Numeric
@@
 Hb Hemoglobin level Numeric
+III-A2 Pre-processing
+Before using the dataset, we preprocessed it to ensure data consistency and reliability.
+Therefore, the goal of pre-processing was to include the highest number of data and have, at the same time, a balance between patients at HF risk and not.
+The following pre-processing steps were performed:
+  • Date values were converted into a standardized date-time format, and a new feature, “lifespan,” was introduced to represent the number of days between the initial characterization date and the recorded outcome date (“death”).
+  • We hold patients for whom we had information on their death or the follow-up period was larger or equal to three years. At the end of this step, 464 patients remained in the dataset.
+  • Patients with at least one missing value in the features used by our model were removed from the dataset.
+At the end of this step, we removed 99 patients from the dataset, and then 365 patients remained.
+  • Patients with inconsistent values or placeholder values in terms of classification (intermediate values) have been removed.
+At the end of this step, we removed 8 patients from the dataset, and we reached the final dimension of the dataset of 357 patients
+  • As for the outcome variable (or also dependent variable), we used a binary classification for the values in the dataset.
+We employed a threshold-based labeling approach, using a three-year threshold to classify patients into two groups: “HF at risk” and “not at HF risk.” The choice of this threshold is justified by technical reasons and clinical bases.
+Indeed, from a technical point of view, this threshold divides the dataset almost perfectly; in fact, we have 45% of samples with the label “at risk” and 55% of samples with the label “not at risk.” This also allowed the predictive model to exhibit good prediction performance of the outcome values.
+From a clinical perspective, the chosen threshold aligns with a time frame during which close monitoring and proactive management are essential to mitigate adverse outcomes in patients with HF.
 III-B Predictive Model Workflow
-Our predictive model is based on an ensemble learning approach, a modified stacking technique.
-Standard stacking combines multiple heterogeneous specialized models trained on the same feature set, our model employs domain-specific feature segregation to enhance predictive accuracy.
-Specifically, the input data are divided into two distinct feature groups: clinical features, including demographic data, patient history, laboratory results, and risk factors, and echocardiographic features.
-Each feature group is processed separately using two identical ML models, each trained exclusively on one of the feature sets.
+Our predictive model (also meta-model or simply model from here onwards) is based on an ensemble learning approach, a modified stacking technique, to classify patients at HF risk or not.
+Standard stacking combines multiple heterogeneous models trained on the same set of features, our approach leverages domain-specific feature segregation to enhance predictive accuracy.
+The input data are divided into two distinct feature groups: clinical features, including demographic data, patient history, laboratory results, and risk factors, and echocardiographic features.
+Each feature group is processed separately using two identical ML models (i.e., also clinical and echocardiographic specialized models from here onwards), each trained exclusively on one of the feature sets.
 This separation ensures that each model specializes in learning patterns from its respective data type without interference from other features.
-In our implementation, both the specialized models and the meta-model are constructed using Random Forest algorithms.
-In detail, each specialized model is tuned using a five-fold cross-validation strategy, where the training data are partitioned into five subsets.
-In detail, for each specialized model, we employ a five-fold cross-validation strategy using GridSearchCV, which systematically tests various combinations of hyperparameters, such as the number of trees, maximum depth, and minimum samples required for a split, to determine the best configuration that minimizes prediction error.
-Once these models are optimized, a single train-test split (with 20% of the data reserved for testing) is utilized to evaluate the final performance.
-The outputs of the two optimized specialized models on the test set are then used as input features for the meta-model.
-Our predictive model workflow is inspired by the research presented by Pinem et al.
+Both the specialized models are implemented using Logistic Regression Classifiers.
+In detail, we employ a five-fold cross-validation strategy using GridSearchCV, which systematically tests various combinations of hyperparameters for a split to determine the best configuration that maximizes the accuracy.
+In addition to these two models, we employ a third model trained on the complete set of features (i.e., the merge between clinical and echocardiographic features).
+Indeed, we used a Random Forest Classifier to capture complex and non-linear relationships across the feature space that could be difficult to detect for Logistic regression classifiers.
+Using all the features together also allows for capturing interactions and patterns that span across different data types, potentially uncovering valuable insights that might be lost when the two sets of features are treated separately.
+A single train-test split (with 20% of the data reserved for testing) is exploited to evaluate the performance of these three models.
+The selection of the three models used was informed by preliminary empirical evidence that indicated their effectiveness alone and in the combined use in the meta-model, given the specific domain of our research and the peculiarities of our dataset, .
+The outputs of the three models described above, along with their confidence in the prediction, are used as inputs for our meta-model (i.e., a classifier), which is implemented as a Logistic Regression Classifier responsible for aggregating the predictions of these three models into a single prediction.
+The idea underlying our predictive model is inspired by the research presented by Pinem et al.
 [7], where the authors trained multiple models on the same set of features to predict heart failure risk.
 In that work, the authors used the output of the models as input for a Voting Classifier.
-In our case, the meta-model on top is also trained on the output of the specialized models, but the main difference is that each specialized model is trained using different features.
-The outcome variable of the two models (i.e., the output of the two specialized models) is the estimated lifespan of a patient.
-The outputs of the two specialized models represent the input for a meta-model, which integrates their predictions to generate the final category for the patients.
-Specifically, the final output of the meta-model is expressed in terms of estimated lifespan.
-As mentioned before, the individuals for whom the model predicts a lifespan of less than 440 days are considered as patients “at risk,” while the other are considered as “not at risk.” In the context of the PrediHealt project, the identification of patients at HF risk allows them to participate in a telemonitoring program, with the objective of acting promptly to let them receive medical attention.
-Integrating two specialized models within a metamodel offers several potential advantages over a single ML model trained on all features. Specifically, these advantages are:
+In contrast, our method introduces two key differences: (i) the use of specialized models trained on disjoint subsets of features (supported by a model trained on the complete feature set) and (ii) the use of a learnable meta-model (rather than a static voting mechanism) for the prediction.
+Integrating a set of models within a meta-model offers several potential advantages over a single ML model:
   1. 1.
-    Improved Interpretability and Modularity.
+    Interpretability and Modularity.
 By separating clinical and echocardiographic features, the contribution of each data type can be analyzed independently.
-This modular approach allows for better insight into how different variables influence predictions.
+Additionally, a third model trained on all features captures joint relationships across the data.
+The study of these models separately also allows for a clearer interpretation of which type of information most influences the final predictions.
+We studied these aspects in our empirical assessment of the proposed solutions.
   2. 2.
-    Reduced Risk of Overfitting.
+    Risk of Overfitting.
 A single model trained on all features may overfit due to redundant or spurious correlations.
@@
   3. 3.
-    Increased Robustness.
+    Robustness.
 This structure makes the system less sensitive to missing data.
@@
     Data availability. We built a model based on variables readily available in clinical reality. All the components used represent part of a standard cardiac evaluation [17].
-  5. 5.
-    Greater Adaptability and Scalability.
-If new predictive techniques emerge for clinical or echocardiographic data, only the respective model needs to be updated, rather than retraining an entire model on all the features.
-This characteristic allows for continuous improvement of the predictions while maintaining the same structure of the predictive framework.
 IV Study Design
@@
 IV-A Research Question
-The primary objective of the experimental assessment is to assess preliminarily the validity of the developed predictive model for the stratification of patients at HR risk.
-In particular, we aimed to show whether the proposed predictive model, which is based on two specialized models leveraging clinical and echocardiographic features and then a meta-model to combine the predictions of these two models, outperforms traditional ML models trained on the entire feature set (i.e., the features used in the two specialized models).
-To that end, we considered three baseline models–i.e., Decision Tree, Random Forest, and Support Vector Regression (SVR)– and compared their results (e.g., accuracy, precision, and sensitivity) with those of our prediction model.
-We choose these baselines because they represent the leading ML techniques in several AI application contexts and medical domains [16], in particular.
-  • Decision Tree.
-It is a tree-structured model that recursively splits the data based on feature values to predict an outcome.
-It is intuitive and interpretable and can handle both classification and regression tasks.
-  • Random Forest.
-It is an ensemble learning method that builds multiple decision trees on random subsets of data and features and aggregates their predictions.
-It improves generalization by reducing overfitting (e.g., as compared with the Decision tree) and increasing robustness.
-It is well-suited for handling high-dimensional data and capturing complex patterns.
-  • Support Vector Regression.
-It is a regression variant of Support Vector Machines that aims to find a function within a specified error margin (epsilon) while maximizing model flatness.
+The primary objective of our initial experimental assessment is to validate the meta-model for the stratification of patients at HF risk.
+In particular, we aimed to show whether it outperforms traditional ML models trained on the entire feature set.
+According to our primary objective, we defined and studied the following main Research Question:
+RQ
+Does our predictive model outperform leading ML techniques in classifying at-risk patients?
+In addition to studying the three ML-based models used in our meta-model, we considered two baselines–i.e., Decision Tree [^3] It is a tree-structured model that recursively splits the data based on feature values to predict an outcome.
+It is intuitive and interpretable and can handle both classification and regression tasks [20].
+and Support Vector Classifier [^4] It is a classification variant of Support Vector Machines [21] designed to identify the optimal hyperplane that separates data points of different classes with the maximum margin.
 It captures non-linear relationships using kernel functions.
-For the baseline models, as we did for our predictive model, the individuals are arranged into the two groups according to the lifespan prediction, i.e., if it is less than 440 days, the individual is considered at risk, otherwise not.
-According to the primary objective of our experimental assessment, we defined and investigated the following main Research Question:
-RQ: Does our predictive model outperform leading ML techniques in classifying at-risk patients?
+(SVC).
+We choose these baselines because they represent the standard for comparison in several AI application contexts and medical domains [22, 23, 24].
 IV-B Experimental Methodology
-To study RQ, we conducted a comparative analysis between our model and the three baselines mentioned before.
-These baselines, differently from our model, were trained using the entire set of 33 selected features in Table I.
-All models were trained and evaluated using the final dataset, consisting of 182 patient records, with 20% of the data reserved for testing.
+To study RQ, we conducted a comparative analysis between our model and the baselines that were trained on the entire set of 33 features in Table I.
+All models considered in the experimental assessment were trained and evaluated using the final dataset with 20% of the data reserved for testing.
 Baselines were implemented using standard algorithms from the scikit-learn library in Python.
@@
 IV-C Metrics to Assess the Predictions of the Models
-The evaluation metrics used are standard and widely accepted in the application of ML techniques within the medical domain [17]. The used metrics are:
-  • Accuracy: It represents the proportion of patients that are correctly identified and it is computed as follows:
-    $\text{Accuracy}=\frac{TP+TN}{TP+TN+FP+FN}\%$
+The evaluation metrics used are standard and widely accepted in the application of ML techniques within the medical domain [25]:
+  • Accuracy: It represents the proportion of patients that are correctly identified, and it is computed as: $\frac{TP+TN}{TP+TN+FP+FN}\%$.
 This metric assumes values between 0% and 100%.
@@
 Instead, FPs (False Positives) represent patients mistakenly identified as at risk, while FNs (False Negatives) represent patients mistakenly identified as not at risk.
-  • Precision (for the at-risk class): It indicates the proportion of patients predicted as at risk who truly are at risk. It is calculated as follows:
-    $\text{Precision}=\frac{TP}{TP+FP}\%$
-    Also, the precision metric assumes values between 0% and 100%. The higher the value, the better the correctness of the prediction is.
-  • Sensitivity (Recall for the at-risk class): It measures the model’s ability to correctly identify patients at risk and is defined as follows:
-    $\text{Sensitivity}=\frac{TP}{TP+FN}\%$
-    Also, this metric assumes values between 0% and 100%. The higher the value, the better the sensitivity of the model is.
-  • F1-Score (for the at-risk class): It is the balanced harmonic mean of precision and sensitivity.
-It offers a balanced measure when the costs of false positives and false negatives differ and is computed as follows:
-    $F1\text{-Score}=2\times\frac{\text{Precision}\times\text{Sensitivity}}{\text{Precision}+\text{Sensitivity}}\%$
-    This metric assumes values between 0% and 100%. The higher the value, the better it is.
-  • Diagnostic Odds Ratio (DOR): It is a comprehensive metric that integrates sensitivity and specificity. It is computed as follows:
-    $DOR=\frac{TP\times TN}{FP\times FN}$
-    The higher the value for DOR, the better it is.
-It is important to highlight that, since the model is intended to stratify patients by identifying those at risk for inclusion in the telemonitoring program, minimizing FN is a primary objective.
-In this context, high precision is particularly desirable, as it ensures that most of the patients flagged as “at risk” truly need intervention.
+  • Precision: It indicates the proportion of patients predicted as at risk who truly are at risk and is calculated as: $\frac{TP}{TP+FP}\%$.
+Also, precision assumes values between 0% and 100%.
+The higher the value, the better it is.
+  • Sensitivity (or recall): It measures the proportion of actual positive patients that are correctly identified as having HF risk.
+It is computed as follows $\frac{TP}{TP+FN}\%$.
+Also, this metric assumes values between 0% and 100%.
+The higher the value, the better it is.
+  • F1-Score: It is the balanced harmonic mean of precision and sensitivity and is computed as: $2\times\frac{\text{Precision}\times\text{Sensitivity}}{\text{Precision}+\text{Sensitivity}}\%$.
+This metric assumes values between 0% and 100%.
+The higher the value, the better it is.
+  • Diagnostic Odds Ratio (DOR): It is a comprehensive metric that integrates sensitivity and specificity.
+It is computed as follows: $\frac{TP\times TN}{FP\times FN}$.
+The higher the value, the better it is.
+It is important to highlight that since the model is intended to stratify patients by identifying those at risk for inclusion in the telemonitoring program, minimizing FN is a primary objective.
+In this context, high precision is particularly desirable, as it ensures that most of the patients at risk truly need to take part in the telemonitoring program.
 Nevertheless, achieving high values across all evaluation metrics remains a desirable goal to ensure overall model reliability.
-TABLE II: Medical evaluation metrics of the proposed model
-Metric Value Explanation of the Results
-Accuracy 84% 84% of patients are correctly identified as at risk.
-Precision (At-Risk) 79% 79% of patients flagged as at risk truly are at risk.
-Sensitivity (At-Risk) 95% 95% of at-risk patients are correctly identified.
-F1-Score (At-Risk) 86% Good balance between precision and sensitivity.
-DOR 45.6 Strong discriminative ability.
+TABLE II: Medical evaluation metrics of the proposed model and specialized ones
+Model Metric Value
+Meta-model Accuracy 78%
+Precision 70%
+Sensitivity 91%
+F1-Score 79%
+DOR 20
+Clinical model Accuracy 72%
+Precision 67%
+Sensitivity 79%
+F1-Score 78%
+DOR 16
+Echocardiographic model Accuracy 78%
+Precision 71%
+Sensitivity 88%
+F1-Score 78%
+DOR 16.31
+Random Forest Classifier Accuracy 76%
+Precision 69%
+Sensitivity 88%
+F1-Score 76%
+DOR 14.5
 V Results
-In this section, we highlight the main findings. We conclude by showing possible practical implications and threats that could affect the validity of the obtained results.
+In this section, we highlight the main findings. We conclude by showing possible practical implications and threats that could affect the validity of the results obtained.
 V-A Main Findings and Discussion
-The performance results of our predictive model are summarized in Table II.
-While the overall accuracy of 84% may appear moderate at first glance, the high sensitivity (95%) ensures that nearly all high-risk patients are correctly identified—an essential requirement for patient monitoring in the PrediHealth project.
-The precision of 79% reflects a relatively high number of false positives, which is acceptable given the clinical context and the overall objective of the PrediHealth project by minimizing FNs.
-Furthermore, the high Diagnostic Odds Ratio (45.6) underscores the model’s strong discriminative power between high- and low-risk patients, reinforcing its suitability for deployment within PrediHealth.
-The results in Table III suggest that the proposed model significantly outperforms the baseline machine learning models.
-It achieves an accuracy of 84%, compared to 57% for the Decision Tree, 60% for the Random Forest, and 62% for SVR.
-Notably, the sensitivity of our model reaches 95% which is well above the 60% (Decision Tree and Random Forest) and 50% (SVR) by suggesting a strong ability to correctly identify high-risk patients as compared with the baseline models.
-The precision of our model (79%) is slightly higher than that of SVR (71%), which has the highest precision among the baseline models.
-This indicates that our model has a better balance between minimizing false positives and ensuring the inclusion of at-risk patients in the telemonitoring program.
-Additionally, the proposed model achieves an F1-score of 86%, outperforming Decision Tree 60% and Random Forest (62%), and Support Vector Regression (59%).
-Its DOR (45.6) further confirms its superior ability to distinguish between high-risk and low-risk patients, compared to Decision Tree (1.69), Random Forest (2.14), and SVR (3.25).
-These improvements highlight the model’s reliability for HF risk stratification of patients.
+The results of our meta-model are summarized in Table II (top).
+Although the overall precision of 78% appears moderate at first glance, the high sensitivity (91%) ensures that nearly all patients at HF risk are correctly identified - an essential requirement for patient monitoring in the PrediHealth project.
+The precision of 70% reflects a relatively high number of false positives, which is acceptable given the overall objective of the PrediHealth project by minimizing FNs.
+Furthermore, the high DOR (20) underscores the model’s strong discriminative power between patients at HF risk and those not.
+This reinforces the suitability of our meta-model for its deployment within the PrediHealth project.
+In Table II, we also report the results for each model underlying our meta-model.
+We can note that the meta-model allows obtaining better results as compared with its underlying ML-based models.
+The only difference concerns the precision value of the Echocardiographic model, which is 71%, while the precision value is 70% for the meta-model.
+These results support the conclusion that combining the predictions of the three underlying models contributes to improved performance, indicating that each model captures complementary aspects of the data.
 TABLE III: Medical evaluation metrics of the baseline ML models
 ML Model Metric Value
-Decision Tree Accuracy 57%
-Precision (At-Risk) 60%
-Sensitivity (At-Risk) 60%
-F1-Score (At-Risk) 60%
-DOR 1.69
-Random Forest Accuracy 60%
-Precision (At-Risk) 63%
-Sensitivity (At-Risk) 60%
-F1-Score (At-Risk) 62 %
-DOR 2.14
-Support Vector Regression Accuracy 62%
-Precision (At-Risk) 71%
-Sensitivity (At-Risk) 50%
-F1-Score (At-Risk) 59%
-DOR 3.25
+Decision Tree Accuracy 67%
+Precision 60%
+Sensitivity 79%
+F1-Score 68%
+DOR 4.81
+SVC Accuracy 68%
+Precision 63%
+Sensitivity 72%
+F1-Score 68%
+DOR 4.76
+The results in Table III suggest that the meta-model outperforms the baseline models.
+It achieves an accuracy of 78%, compared to 67% for the Decision Tree and 68% for SVC.
+The sensitivity of our model reaches 91% which is well above the 79% and 72% for Decision Tree and SVC, respectively.
+This suggests a better capability of the meta-model to correctly identify patients at HF risk and not if compared with the baselines.
+The precision of the meta-model (70%) is higher than the precision of the baselines (60% and 63% for Decision Tree and SVC, respectively)..
+This indicates that the meta-model has a better balance between minimizing FPs and ensuring the inclusion of at-risk patients in the telemonitoring program of PrediHealth.
+Additionally, the proposed model achieves an F1-score of 79%, outperforming Random Forest and SVC (68% for both models).
+Its DOR (20) further confirms its superior ability to distinguish between patients at HF risk and those not, compared to Decision Tree (4.81) and SVC (4.76).
+These improvements highlight the reliability of the model for the aim for which it is developed.
 V-B Implications
@@
 This choice reflects our view that the project’s implications are already evident, and it allows us to emphasize that our findings hold significance beyond the boundaries of the project itself.
-Specifically, we developed a predictive model for HF risk stratification of patients using a modified stacking ensemble approach applied to clinical and echocardiographic data.
-It achieved high sensitivity while maintaining an adequate accuracy, using a 440-day lifespan threshold.
-This outcome suggests that a modified stacking ensemble approach can effectively identify a significant portion of high-risk HF patients.
-This is relevant for practitioners in the healthcare field, as the model functions as a valuable screening tool.
-Its high sensitivity supports its role in initial patient stratification, prompting closer monitoring.
-Researchers could be interested in improving our model in order to enhance its accuracy while maintaining high sensitivity, potentially by integrating additional data modalities or optimizing our stacking ensemble approach.
-Our overall model, which separates clinical and echocardiographic features into distinct specialized models before integrating their predictions via a meta-model, presents a potential advantage, as our initial results suggest.
-Similar to how modularity improves software systems, this structure may enhance the interpretability of risk scores for clinicians.
-Practitioners could benefit from this increased transparency, as it allows them to discern whether a patient’s risk is primarily driven by general clinical status or specific cardiac parameters derived from echocardiography.
-Researchers should empirically validate the interpretability and robustness of this architecture, comparing it to monolithic models trained on all features flatly.
-Our initial results seem to justify this future research direction and also to explore optimal meta-learning strategies for integrating predictions from specialized models by assessing the scalability when incorporating additional features.
-The predictive model was trained and evaluated on real patient data/records, which underwent extensive pre-processing, including feature selection and the removal of records with missing data, ultimately resulting in a dataset of 182 instances.
-Researchers could validate the model’s performance and generalizability on larger, multi-center datasets representing diverse patient populations.
+Specifically, we developed a predictive model using a modified stacking ensemble approach applied to clinical and echocardiographic data.
+It achieved high sensitivity while maintaining adequate accuracy.
+This outcome suggests that a modified stacking ensemble approach can effectively identify a significant portion of patients at HF risk.
+This is relevant for practitioners in the healthcare field, as the model works as a valuable screening tool.
+Its high sensitivity supports its role in initial patient stratification.
+Researchers could be interested in improving our model to enhance its accuracy while maintaining high sensitivity, potentially by integrating additional data modalities or optimizing the used stacking ensemble approach.
+Our meta-model, which uses clinical and echocardiographic features and all the features flattened, presents a potential advantage, as our initial results suggest.
+Similar to how modularity improves software systems, this structure may enhance the interpretability of risk scores for practitioners.
+They could benefit from this increased transparency, as it allows them to discern whether a patient’s risk is primarily driven by general clinical status or specific cardiac parameters derived from echocardiography.
+This aspect is also relevant for researchers.
+They could also be interested in empirically validating the interpretability and robustness of this architecture.
+For example, we provided initial evidence by showing that combining the predictions of three ML-based models by a meta-model improves the final predictions.
+Therefore, our empirical assessment seems to justify future research on the use of the ensemble learning approach and to explore optimal meta-learning strategies for integrating predictions from different models.
+The predictive model was trained and evaluated on real patient data/records, which underwent pre-processing, including feature selection and the removal of records with missing data.
+Researchers could validate the performance and generalizability of the model on larger multi-center datasets representing diverse patient populations.
 A particular focus on HFpEF cases, as discussed in our related work, would be valuable.
-Researchers could also be interested in investigating the impact of different missing data imputation techniques and the development of automated, context-aware feature selection methods to enhance the model’s robustness and reliability.
+Researchers could also be interested in investigating the impact of different missing data imputation techniques and the development of automated, context-aware feature selection methods to enhance the robustness and reliability of a predictive model.
 V-C Threats to Validity
@@
 V-C1 External validity
-Our predictive model was developed and evaluated using 182 samples.
+Our predictive model was developed and evaluated using 357 samples.
 Generalizing outcomes from this clinically specific cohort poses a threat to external validity.
@@
 V-C2 Internal Validity
-Internal validity threats mostly concern factors internal to our study design that might have influenced the results.
-The main internal validity threat of the study concerns the limited number of baselines used for comparison.
-This is because this does not exclude that there could be other ML models with better performance.
+The main internal validity threat of the study concerns the number of baselines used in the empirical assessment.
+There could be other ML models outperforming our meta-model.
 On the other hand, the models used as comparison baselines are among the most used in the medical field, justifying their choice.
 V-C3 Construct Validity
-As for this kind of threats to validity, one key concern is how we define “HF risk” as survival of fewer than 440 days after characterization.
-While this provides a clear target, it may not fully capture the broader clinical view of when a patient needs intervention or faces a high risk of worsening health, which can lead to serious complications beyond death.
-As for the metrics used to assess our predictive model and baselines, we used standard performance measures such as sensitivity and accuracy.
+As for this kind of threat to validity, one key concern is how we define patients at HF risk.
+It may not fully capture the broader clinical view of when a patient needs intervention or faces a high risk of worsening health, which can lead to serious complications beyond death.
+As for the metrics used to assess our predictive model and baselines, we used standard performance measures.
 While these provide useful information, they possibly do not give a complete picture.
-Lastly, the model assumes that medical records are accurate.
-Any errors in these records could reduce the reliability of predictions and limit its usefulness in real-world settings.
+Lastly, our model assumes that medical records are accurate.
+Any errors in these records could reduce the reliability of predictions and limit their usefulness in real-world settings.
 V-C4 Conclusion validity
-Our definition of the binary outcome (“at risk” vs.
-“not at risk”), chosen for technical reasons (dataset balance) and clinical relevance, could have affected outcomes.
+Our definition of the binary outcome (at risk vs.
+not at risk), chosen for technical reasons (dataset balance) and clinical relevance, could have affected outcomes.
 Using a different threshold could lead to different conclusions.
 Another internal validity threat concerns the choice of 33 features, which was based on clinical experience and data availability.
-This selection might leave out important predictors or include less useful ones, making it harder to fully represent the complex nature of HF risk.
+This might leave out important features or include less useful ones.
 V-C5 Reliability validity
@@
 This inherently limits direct replication.
-To mitigate this, we have detailed the features used in the model and the design of the empirical study to assess its validity.
-VI Conclusions
-In this paper, we propose a predictive model based on machine learning (ML) techniques to identify patients at risk of HF.
-The model adopts an ensemble learning approach using a modified stacking technique, which combines the outputs of two specialized models—one based on clinical features and the other on echocardiographic features—via a meta-model.
-We evaluate the model on a real-world dataset, and the results demonstrate strong performance in stratifying patients by HF risk.
-In particular, the model achieves high sensitivity (95%), ensuring that nearly all high-risk patients are correctly identified.
-The overall accuracy is 84%, which, while moderate in general ML contexts, is acceptable given our priority of minimizing false negatives (patients mistakenly identified as at risk).
-This is especially relevant for the telemonitoring program of the PrediHealth research project.
-Indeed, our preliminary findings also suggest that ML-based risk stratification models can serve as valuable decision-support tools—not only within the PrediHealth initiative but also more broadly in clinical practice—to facilitate early intervention and personalized patient management.
-We also compared our model against three baselines that treat features in a flat manner (i.e., without grouping them into clinical and echocardiographic categories).
-The results indicate that our model consistently outperforms these baselines, highlighting the benefit of feature specialization and structured integration.
-Future research will focus on expanding the dataset to further validate the model, optimizing its interpretability for clinical applications, and integrating additional physiological and behavioral parameters to improve prediction.
-It is also worth noting that the current model was trained on real-world patient data from the Tuscany region of Italy.
-Due to the limited dataset size and its regional scope, the findings should be interpreted with caution; however, the promising initial results motivate conducting broader studies.
+We aim to make the source code implementing the data pre-processing pipeline and the described ensemble model architecture available upon reasonable request, facilitating conceptual replication and verification of our research.
+VI Conclusion
+In this paper, we propose a predictive model based on machine learning (ML) techniques to identify patients at HF risk.
+The model adopts an ensemble learning approach using a modified stacking technique, which combines the outputs of three underlying ML-based models.
+We evaluate our model on a real-world dataset, and the results demonstrate strong performance in stratifying patients at HF risk.
+In particular, the model achieves high sensitivity (91%), ensuring that nearly all patients are correctly identified as being at HF risk or not.
+The overall accuracy is 78%, which is acceptable given our priority of minimizing false negatives (patients mistakenly identified as at risk).
+This is especially relevant for the telemonitoring program of the research project PrediHealth, for which the proposed predictive model has been developed.
+We also compared our model against some ML-based models, and the results indicate that our model outperforms these models.
 Acknowledgment
 This project has been financially supported by the European Union NEXTGenerationEU project and by the Italian Ministry of the University and Research MUR (M4C2) as part of the project “Tuscany Health Ecosystem” - THE - Spoke 10 - CUP - J13C22000420001.
+This work was also partially supported by the project “DHEAL – COM- Digital Health Solutions in Community Medicine” under the Innovative Health Ecosystem (PNC) - National Recovery and Resilience Plan (NRRP) program funded by the Italian Ministry of Health.